OpenStudy (anonymous):

how about this one. solve for x by factoring. x squared -3x=4. show work please :)

OpenStudy (lgbasallote):

x^2 -3x = 4? x(x-3) = 4 equate to 4 each factor x = 4; x-3 = 4 so..x=4; x = 7

OpenStudy (accessdenied):

i dont think you can necessarily do that x(x-3) = 4 does not necessarily imply that x and x-3 would = 4

Directrix (directrix):

x^2 -3x - 4 = 0 (x - 4)(x + 1) = 0 By the Zero Product Property x - 4 = 0 or x+1 = 0 x = 4 or x = -1.

OpenStudy (accessdenied):

you could reasonably say that "x and x-3 are factors of 4" in which case, x=4 and x-3 =4-1 = 1 is a reasonable factor is true but then you need another pair of x and x-3 that multiply together to get 4

Directrix (directrix):

There is NO Four Product Property. ======================== x(x-3) = 4 does NOT imply that x = 4 and (x-3) = 4 or any other values for that matter.

OpenStudy (accessdenied):

x=-1 and x-3 = -1-3 = -4 is also a true pair of factors for 4 -1*-4 = 4 although i dont think its as easy in all cases to use this kind of method, just sayin'. D: